Determine the total hardness as CaCO3 of a sample of water that contains 0.81 mg/L of Ca(HCO3)2 and 0.73 mg/L of Mg(HCO3)2. Equivalent weights of Ca(HCO3)2 and Mg(HCO3)2 are 81.10 g/mol and 73.17 g/mol, respectively.


1
Expert's answer
2021-04-19T04:34:18-0400

"ppm (CaCO_3) = \frac{( \frac{0.73}{146}+ \frac{0.73}{162})×100}{100}×10^6=10^4ppm"